# Seeing Beyond Haze: Generative Nighttime Image Dehazing

### v1

@@ 4 Experiments
In our experiments, we test on a real-world nighttime dataset and use five no-reference image quality assessment metrics, namely MUSIQ [23], ManIQA [57], ClipIQA [51], HyperIQA [48], and TRES [12], to calculate the quality scores of the dehazed results.
RealNightHaze is a real-world nighttime haze dataset consisting of 440 images. These images are sourced from previous studies [65, 21, 31]. Due to the complex degradation factors in RealNightHaze, including haze, glow effects, extreme darkness, and noise, this dataset provides a comprehensive basis for evaluating the performance of existing methods.
@@ 4.1 Implementation Details
Generative Dehazing Priors Our generative dehazing priors are obtained from a pre-trained dehazing model and a pre-trained diffusion model. In this paper, the dehazing model is a transformer network [16] pre-trained with strong augmentations [31], while Stable Diffusion v2 serves as the large-scale pre-trained diffusion model. To preserve the generative capability of the pre-trained diffusion model, we integrate a LoRA module [45]. The rank of the low-rank decomposition in LoRA is set to 8, with a 10% dropout applied to the LoRA layers. Additionally, the LoRA $\alpha$ parameter is set to 16 to control the scaling of the learned updates.
During training, we utilize the dehazing model to construct training pairs from real-world data. These pairs along with their text prompt are then used to fine-tune the diffusion model. The training batch size is set to 12, with a gradient accumulation step of 4 for stable training. The learning rate is set to $2\times 10^{-4}$ for the U-Net and $4\times 10^{-5}$ for the text encoder, with an input resolution of $512\times 512$. The total number of training steps is set to 10,000.
Figure 3: Qualitative results on the real-world haze dataset from NightDeFog’20 [56], NightEnhance’23 [21], SFSNiD’24 [7], and our method. Our method not only suppresses nighttime haze effects but also infers missing backgrounds in severely degraded regions.
Controllable Generative Dehazing The generative capabilities of our diffusion model are enhanced by using training pairs from the detail enhancement and severe degradation models. During training, we utilize the detail enhancement model [20] to recover fine-scale details in the initial dehazed images obtained from the pre-trained dehazing model, resulting in the detail enhancement image pairs. Additionally, we adopt the degradation strategy described in Section 3.2 to generate training pairs for background synthesis, where the severity ratio $S$ is approximately 98%. These paired datasets, along with their respective text prompts, are used to train our diffusion model. The customized text prompt for initial dehazing pairs $t_{\rm{ID}}$ is “A dehazed image with slight degradation.”, while the text prompt for detail enhancement pairs $t_{\rm{DE}}$ is “A high-detail, dehazed image with slight degradation.”. The prompt for background synthesis pairs $t_{\rm{bs}}$ is “An image with no degradation, generation.”. Since the only difference between the text prompts for initial-dehazing and detail enhancement pairs is the term ”high-detail,” this distinction can be effectively learned by our diffusion model.
In each training batch, we sample paired data equally from the initial-dehazing, detail-enhancement, and background-synthesis pairs to ensure effective training. During inference, given a real-world nighttime haze image, we utilize the fine-tuned diffusion model to enhance its visibility. With the text prompt “A high-detail, dehazed image with no degradation, generation”, our fine-tuned diffusion model produces highly generative dehazed results with synthesis of details and background content. The text prompt ”A dehazed image with slight degradation” would instead lead to low-generative results that do not contain synthesized details or backgrounds, yielding greater factual accuracy but lower visual quality.
@@ 4.2 Quantitative Evaluation on RealNightHaze
We evaluate the performance of our BeyondHaze method on the RealNightHaze dataset, with experimental results presented in Table 1. We compare our method against several state-of-the-art techniques, including Uformer [54], Restormer [61], DiT [39], NightDeFog [56], NightEnhance [21], and SFSNiD [7]. Uformer, Restormer, and DiT are restoration backbones trained on synthetic datasets, while the remaining methods are specifically designed for nighttime image dehazing.
Table 1 shows that our BeyondHaze achieves notable improvements across all non-reference metrics. For models trained on synthetic datasets, such as Restormer and DiT, their MUSIQ scores remain around 45. In contrast, existing nighttime image dehazing methods, such as NightDeFog and NightEnhance, demonstrate greater effectiveness in enhancing visibility in hazy images. For example, NightDeFog achieves MUSIQ and TRES scores of 51.61 and 68.88, respectively. By comparison, our method achieves MUSIQ and TRES scores of 65.79 and 80.08, surpassing NightDeFog by 14.18 and 11.2 points, respectively.
Figure 4: Qualitative results from NightDeFog’20 [56], NightEnhance’23 [21], SFSNiD’24 [7], and our method on real-world datasets. “Ours-low” refers to our low-generative results, which prioritize haze removal, while “Ours-high” refers to our high-generative results, allowing the network to infer details and backgrounds in severely degraded regions. The first two rows exhibit the ability of Ours-high to infer missing scene content, while the last two rows display fine-scale detail generation by Ours-high. Zoom in for better visualization.
@@ 4.3 Qualitative Evaluation on RealNightHaze
Figure 3 presents a qualitative comparison of various methods on real-world nighttime haze data. It can be observed that existing nighttime image dehazing methods often suffer from noise artifacts and tend to produce dark outputs. In contrast, our method effectively suppresses haze effects, enhancing overall image visibility. Moreover, we find that existing nighttime image dehazing methods struggle to infer backgrounds in severely degraded regions, such as those affected by dense haze or strong glow. As shown in Figure 3 and Figure 4, these methods tend to generate black predictions in such regions. The primary reason is that information in these regions may be absent. Since these methods lack generative capabilities, they are unable to infer backgrounds, leading to inferior performance. In contrast, our BeyondHaze leverages strong priors from a pre-trained diffusion model to not only address haze effects but also infer backgrounds in regions entirely occluded by dense haze. As shown in Figure 3 and Figure 4, our method achieves significant performance improvements.
@@ 4.4 Ablation Studies
Analysis of the Generative Dehazing Priors Our pre-trained dehazing model achieves a MUSIQ score of 58.92, a TRES score of 63.83, and a ClipIQA score of 0.5290. In contrast, our finetuned diffusion model achieves a MUSIQ score of 60.29 and a ClipIQA score of 0.6217, surpassing the pre-trained dehazing model by 1.37 in the MUSIQ score. This demonstrates that distilling knowledge from a pre-trained dehazing model to an image diffusion model leads to higher performance than the dehazing teacher model.
GDP Supplementary Models MUSIQ TRES ClipIQA
@@
Table 2: Ablation studies on real-world haze datasets. “GDP” denotes our generative dehazing priors. We use three non-reference metrics for evaluation: MUSIQ [23], TRES [12], and ClipIQA [51]. The scores in the first row are calculated from the results produced by our pre-trained dehazing model.
Analysis of Controllable Generative Dehazing Table 2 shows that our supplementary models designed for generative details and backgrounds lead to a substantial performance improvement. Including them yields a MUSIQ score of 65.79, outperforming the pre-trained dehazing model by 6.87 points. This improvement demonstrates that the two models are effective at elevating the generative capabilities of the pre-trained diffusion model, resulting in superior performance.
Analysis of Different Generative Levels Using text prompts, our fine-tuned diffusion model can control the generative level of the outputs. The experimental results are shown in Figure 4. Our low-generative results, which prioritize haze removal, effectively mitigate haze effects but fail to infer background details and content in severely degraded regions. In contrast, our high-generative results enable the network to infer background details and content in these regions, significantly boosting visual quality.
@@ 5 Conclusion
This paper presents BeyondHaze, a novel generative approach for nighttime image dehazing that integrates dehazing priors with generative capabilities. We first pretrain a dehazing model using strong augmentations that simulate noise and light effects in hazy night scenes, allowing the model to learn strong dehazing priors. This knowledge is then distilled into a diffusion model in a manner that preserves its generative capabilities. The generative power of the diffusion model for recovering fine-scale details and missing background regions in nighttime haze images is strengthened through training data from two additional models: a detail enhancement model and a severe degradation model. The detail enhancement model refines dehazed regions with lost details to produce detail enhancement training pairs, while the severe degradation model creates training pairs from clear images with severe augmentations to aid the diffusion model in learning to synthesizes missing background areas. Customized text prompts are associated with these training pairs to allow users to control the level of generative dehazing in the outputs. Extensive experimental results on real-world nighttime datasets demonstrate that BeyondHaze achieves substantial improvements over other methods.
Supplementary Material
@@
Figure S1: Qualitative results from DiT’23 [39], NightDeFog’20 [56], NightEnhance’23 [21], SFSNiD’24 [7], and our method, on the real-world dataset. Our method not only reduces dense haze and strong glow effects but also infers missing backgrounds in severely degraded regions. Zoom in for better visualization.
A More Results on Real-world Datasets
@@
In this section, we present additional real-world dehazing results comparing our BeyondHaze with state-of-the-art (SOTA) methods, including DiT’23 [39], NightDeFog’20 [56], NightEnhance’23 [21], and SFSNiD’24 [7]. DiT’23 serves as a general image restoration backbone trained on synthetic datasets, while the other SOTA methods are specialized nighttime image dehazing approaches trained using both labeled synthetic datasets and unlabeled real-world datasets. The experimental results are shown in Figs. S1, S2, S3, S4, S5, S6, S7 and S8. We can find that our BeyondHaze not only reduces dense haze and strong glow effects but also infers missing backgrounds in severely degraded regions.
B Experimental Details
@@
Our BeyondHaze is built upon on generative dehazing priors and controllable generative dehazing. We provide more training details as follow.
B.1 Generative Dehazing Priors
@@
Our pre-trained dehazing model is a ViT-Large [16] architecture with 24 encoder layers and 8 decoder layers. During training, we initialize the model with pre-trained parameters from [16] and fine-tune it using augmentations from [31], where clear images are intentionally degraded with light effects and noise. This approach forces the model to learn robust background priors, enabling it to infer clear scene information despite degraded inputs.
@@
During training, input images are randomly cropped to 224 × 224. We set the total training steps to 20,000 and use a batch size of 128. The Adam optimizer is used with an initial learning rate of 1.5e-4. During inference, we apply overlapping sampling to reconstruct high-resolution images, as ViT-Large only supports an input size of 224 × 224. Specifically, a 224 × 224 sliding window is used to extract overlapping regions from high-resolution hazy inputs. The pre-trained dehazing model is then applied to restore these regions. Since each pixel receives multiple predictions, we average them to obtain the final dehazing results.
B.2 Controllable Generative Dehazing
@@
We utilize the detail enhancement model [20] to recover fine-scale details in the initial dehazed images obtained from the pre-trained dehazing model. This model is built on FLUX.1 [20] and incorporates ControlNet, trained using the augmentation method from [53].
@@
Our severe degradation model severely degrades clear nighttime images by blending light effects and adding noise, creating a heavily augmented dataset, formulated as in [31]:
@@
We use real-world light maps for the augmentation, which is sourced from [31]. At each training step, we randomly select a light map for each clear image. Then, we select various regions and amplify their brightness using Gaussian kernels. The number of selected regions is between 1 and 10, and the size of Gaussian kernels ranges from 15 to 160.
@@
With these parameters, the severity $S$ averages approximately 98% over all training steps. Note that the purpose of our severe augmentation model is to heavily degrade clear nighttime images, simulating real-world scenarios where background details are significantly degraded or entirely lost. It is not intended to perfectly simulate glow/haze effects or transmission maps. Therefore, Gaussian noise, the blend weight map, and the light map are sufficient as long as the augmentation remains severe.

### v2

@@ 4 Experiments
In our experiments, we test on a real-world nighttime dataset and use five no-reference image quality assessment metrics, namely MUSIQ [ke2021musiq], ManIQA [yang2022maniqa], ClipIQA [wang2023exploring], HyperIQA [su2020blindly], and TRES [golestaneh2022no], to calculate the quality scores of the dehazed results.
RealNightHaze is a real-world nighttime haze dataset consisting of 440 images. These images are sourced from previous studies [zhang2020nighttime, lin2024nighthaze]. Due to the complex degradation factors in RealNightHaze, including haze, glow effects, extreme darkness, and noise, this dataset provides a comprehensive basis for evaluating the performance of existing methods.
@@ 4.1 Implementation Details
Generative Dehazing Priors Our generative dehazing priors are obtained from a pre-trained dehazing model and a pre-trained diffusion model. In this paper, the dehazing model is a transformer network [he2022masked] pre-trained with strong augmentations [lin2024nighthaze], while Stable Diffusion v2 serves as the pre-trained diffusion model. To preserve the generative capability of the pre-trained diffusion model, we integrate a LoRA module [ruiz2023dreambooth]. The rank of LoRA is set to 8.
Figure 3: Qualitative results from NightEnhance’23 [jin2023enhancing], SFSNiD’24 [cong2024semi], and our method on real-world datasets. “Ours-low” refers to our low-generative results, which prioritize haze removal, while “Ours-high” refers to our high-generative results, allowing the network to infer details and backgrounds in severely degraded regions. “Gen. Level Map” shows the generative-level maps of “Ours-high,” with red indicating high and blue indicating low generative regions. Zoom in for better visualization.
Figure 4: Qualitative results from NightDeFog’20 [yan2020nighttime], NightEnhance’23 [jin2023enhancing], SFSNiD’24 [cong2024semi], and our method on real-world datasets. “Ours-low” refers to our low-generative results, which prioritize haze removal, while “Ours-high” refers to our high-generative results, allowing the network to infer details and backgrounds in severely degraded regions. The first two rows exhibit exhibits the ability of “Ours-high” to infer missing scene content, while the last two rows display fine-scale detail generation by Ours-high.
Controllable Generative Dehazing The generative capabilities of our diffusion model are enhanced by using training pairs from the detail enhancement [flux1] and severe degradation models. Additional implementation and training details are provided in the supplementary material. During inference, given a real-world nighttime haze image, the text prompts $t_{\text{low}}$ and $t_{\text{high}}$ enable users to control the generative level of the outputs. By processing both prompts, our model also derives a generative-level map that highlights hallucination-prone regions in the high-generative outputs. Further details, including prompt design and inference settings, are provided in the supplementary material.
@@ 4.2 Quantitative Evaluation on RealNightHaze
We evaluate the performance of our BeyondHaze method on the RealNightHaze dataset, with experimental results presented in Table 1. We compare our method against several state-of-the-art techniques, including Uformer [wang2022uformer], Restormer [zamir2022restormer], DiT [Peebles2023DiT], RealFill [tang2024realfill], NightDeFog [yan2020nighttime], NightEnhance [jin2023enhancing], Diff-Dehazer [lan2025exploiting] and SFSNiD [cong2024semi]. Uformer, Restormer, DiT and RealFill are restoration backbones trained on synthetic datasets, while the remaining methods are specifically designed for image dehazing.
Table 1 shows that BeyondHaze consistently improves all non-reference metrics. Synthetic-trained models such as Restormer and DiT reach only  45 MUSIQ, while nighttime dehazing methods like NightDeFog and NightEnhance perform better (e.g., NightDeFog: 51.61 MUSIQ, 68.88 TRES). Our method further increases these scores to 65.79 MUSIQ and 80.08 TRES.
@@ 4.3 Qualitative Evaluation on RealNightHaze
Figure 3 presents a qualitative comparison of different methods on real-world nighttime haze data. Existing nighttime dehazing approaches often suffer from noise artifacts, produce dark outputs, and fail to recover background content in severely degraded regions (e.g., dense haze or strong glow). As shown in Figures 3 and 4, they tend to generate black areas in such cases. The main reason is their lack of generative capability, which prevents them from inferring the missing background content. In contrast, our BeyondHaze leverages strong priors from a pre-trained diffusion model to not only address haze effects but also infer backgrounds in regions entirely occluded by dense haze. As shown in Figures 3 and 4, our method achieves significant performance improvements. Moreover, we introduce a generative-level map during inference to visualize regions prone to generative synthesis or hallucination. Fig. 3 shows that the red-highlighted areas in this map indicate regions likely to contain hallucinated content, such as the house color and texture, building boundaries, and sky regions.
@@ 4.4 Ablation Studies
Generative Dehazing Priors Our pre-trained dehazing model achieves a MUSIQ score of 58.92, a TRES score of 63.83, and a ClipIQA score of 0.5290. In contrast, our finetuned diffusion model achieves a MUSIQ score of 60.29 and a ClipIQA score of 0.6217, clearly surpassing the pre-trained dehazing model. This demonstrates that distilling knowledge from a pre-trained dehazing model to an image diffusion model leads to higher performance than the dehazing teacher model.
Table 2: Ablation studies on real-world haze datasets. “GDP” denotes our generative dehazing priors. We use three non-reference metrics for evaluation: MUSIQ [ke2021musiq], TRES [golestaneh2022no], and ClipIQA [wang2023exploring]. The scores in the first row are calculated from the results produced by our pre-trained dehazing model.
GDP Supplementary Models MUSIQ $\uparrow$ TRES $\uparrow$ ClipIQA $\uparrow$
@@
Controllable Generative Dehazing Table 2 shows that our supplementary models designed for generative details and backgrounds lead to a substantial performance improvement. Including them yields a MUSIQ score of 65.79, outperforming the pre-trained dehazing model by 6.87 points. This improvement demonstrates that the two models are effective at elevating the generative capabilities of the pre-trained diffusion model, resulting in superior performance.
Different Generative Levels Using text prompts, our fine-tuned diffusion model can control the generative level of the outputs. The experimental results are shown in Figure 4. Our low-generative results, which prioritize haze removal, effectively mitigate haze effects but fail to infer background details and content in severely degraded regions. In contrast, our high-generative results enable the network to infer background details and content in these regions, significantly boosting visual quality.
@@ 5 Conclusion
This paper presents BeyondHaze, a novel generative approach for nighttime image dehazing that integrates dehazing priors with generative capabilities. We first pre-train a dehazing model using strong noise and glow augmentations to learn robust priors, which are then distilled into a diffusion model while preserving its generative power. The generative power of the diffusion model to recover fine-scale details and missing background regions in nighttime haze images is enhanced by training data from two auxiliary models: a detail enhancement model and a severe degradation model. The detail enhancement model refines dehazed regions with lost details to produce detail enhancement training pairs, while the severe degradation model creates training pairs from clear images with severe augmentations to aid the diffusion model in learning to synthesizes missing background areas. Customized text prompts are associated with these training pairs to allow users to control the level of generative dehazing in the outputs. A generative-level map is further derived during inference to indicate regions that rely more on generative synthesis than on factual restoration. Extensive experiments on real-world nighttime datasets demonstrate that BeyondHaze achieves substantial improvements over other methods.
@@
Figure 5: Qualitative results from NightDeFog’20 [yan2020nighttime], NightEnhance’23 [jin2023enhancing], SFSNiD’24 [cong2024semi] and our method on real-world data. Our method not only reduces dense haze and strong glow but also infers missing background details and content in severely degraded regions.
Appendix A More Results on Real-world Datasets
@@
In this section, we present additional real-world dehazing results comparing our BeyondHaze with state-of-the-art (SOTA) methods, including DiT’23 [Peebles2023DiT], NightDeFog’20 [yan2020nighttime], NightEnhance’23 [jin2023enhancing], and SFSNiD’24 [cong2024semi]. DiT’23 serves as a general image restoration backbone trained on synthetic datasets, while the other SOTA methods are specialized nighttime image dehazing approaches trained using both labeled synthetic datasets and unlabeled real-world datasets. The experimental results are shown in Figs. 5, 8, 9, 10, 11, 12, 13, and 14. We can find that our BeyondHaze not only reduces dense haze and strong glow effects but also infers missing backgrounds in severely degraded regions.
Figure 6: Qualitative results from NightEnhance’23 [jin2023enhancing], SFSNiD’24 [cong2024semi], and our method on real-world datasets. “Ours-low” refers to our low-generative results, which prioritize haze removal, while “Ours-high” refers to our high-generative results, allowing the network to infer details and backgrounds in severely degraded regions. “Gen. Level Map” shows the generative-level maps of “Ours-high,” with red indicating high and blue indicating low generative regions. Zoom in for better visualization.
Appendix B More Results with Generative Level Maps
In this section, we provide more visualizations of our generative level maps in Figure 6. The red-highlighted areas in the map indicate regions likely to contain hallucinated content, such as the details of trees and the color and brightness of sky regions.
Appendix C Experimental Details
@@
Our BeyondHaze framework is built upon generative dehazing priors and controllable generative dehazing. We provide additional training details as follows.
C.1 Generative Dehazing Priors
Pre-trained Dehazing Model Details
@@ C.1 Generative Dehazing Priors
Our pre-trained dehazing model is a ViT-Large [he2022masked] architecture with 24 encoder layers and 8 decoder layers. During training, we initialize the model with pre-trained parameters from [he2022masked] and fine-tune it using strong augmentations from [lin2024nighthaze], where clear images are intentionally degraded with light effects and noise. This approach forces the model to learn robust background priors, enabling it to infer clear scene information despite degraded inputs.
@@
During training, input images are randomly cropped to 224 × 224. We set the total number of training steps to 20,000 and use a batch size of 128. The Adam optimizer is used with an initial learning rate of 1.5e-4. During inference, we apply overlapping sampling to reconstruct high-resolution images, as ViT-Large only supports an input size of 224 × 224. Specifically, a 224 × 224 sliding window is used to extract overlapping regions from high-resolution hazy inputs. The pre-trained dehazing model is then applied to restore these regions. Since each pixel receives multiple predictions, we average them to obtain the final dehazing results.
Details of Initial Dehazing Pairs
Once the pre-trained dehazing model is obtained, we utilize it to construct training pairs from real-world data. These pairs, along with their text prompts, are then used to fine-tune the diffusion model.
Figure 7: Examples of our severely augmented training pairs. For each triplet, we show the ground-truth image (left), the augmented light map (middle), and the augmented image (right).
C.2 Controllable Generative Dehazing
C.2.1 Detail Enhancement Model
@@
During training, we use the detail enhancement model [flux1] to recover fine-scale details in the initial dehazed images produced by the pre-trained dehazing model, resulting in detail-enhanced image pairs. This model is built on FLUX.1 [flux1] and incorporates a ControlNet module trained using the augmentation strategy from [wang2021real].
C.2.2 Severe Degradation Model
@@
Our severe degradation model heavily degrades clear nighttime images by blending light effects and injecting noise, producing a strongly augmented dataset formulated as in [lin2024nighthaze]:
@@
We use real-world light maps for the augmentation, which is sourced from [lin2024nighthaze]. At each training step, we randomly select a light map for each clear image. Then, we select various regions and amplify their brightness using Gaussian kernels. The number of selected regions is between 1 and 10, and the size of Gaussian kernels ranges from 15 to 160.
@@
With these parameters, the severity $S$ averages approximately 98% over all training steps. Note that the purpose of our severe augmentation model is to heavily degrade clear nighttime images, simulating real-world scenarios where background details are significantly degraded or entirely lost. It is not intended to perfectly simulate glow/haze effects or transmission maps. Therefore, Gaussian noise, the blend weight map, and the light map are sufficient as long as the augmentation remains severe. Figure 7 visualizes examples of the severely augmented training pairs.
C.2.3 Training Details
The customized text prompt for initial dehazing pairs $t_{\rm{ID}}$ is “A dehazed image with slight degradation.”, while the text prompt for detail enhancement pairs $t_{\rm{DE}}$ is “A high-resolution, dehazed image with slight degradation.”. The prompt for background synthesis pairs $t_{\rm{bs}}$ is “An image with no degradation, generation.”. Since the only difference between the text prompts for initial-dehazing and detail enhancement pairs is the term ”high-resolution,” this distinction can be effectively learned by our diffusion model.
In each training batch, we sample paired data equally from the initial-dehazing, detail-enhancement, and background-synthesis subsets to ensure balanced learning. The total training batch size is set to 12, with a gradient accumulation step of 4 for stable training. The learning rate is set to $2\times 10^{-4}$ for the U-Net, with an input resolution of $512\times 512$. The total number of training steps is 10,000. All experiments are conducted using 4 NVIDIA GPUs, each with 24 GB of memory.
During inference, given a real-world nighttime haze image, we utilize the fine-tuned diffusion model to enhance its visibility. With the high-generative prompt $t_{\text{high}}$, “A high-resolution, dehazed image with no degradation, generation”, the model produces highly generative dehazed results with synthesized details and background content. Conversely, the low-generative prompt $t_{\text{low}}$, “A dehazed image with slight degradation”, suppresses generative synthesis, yielding results with higher fidelity but relatively lower visual quality.
Appendix D Ablation Studies
In this section, we present visual ablation results to validate the effectiveness of our generative dehazing priors and the supplementary components. The comparisons are shown in Figure 15. While our pre-trained model learns strong background priors, its outputs still exhibit noticeable artifacts, especially in regions with dense haze or intense glow. In contrast, Ours w/o supp equips the diffusion model with both dehazing and generative capabilities, resulting in more realistic outputs with enhanced visibility and improved background recovery. Finally, Ours, which incorporates both generative priors and supplementary training pairs, further enhances fine-scale details and restores severely obscured scene content. These results highlight the benefits of guided diffusion training and demonstrate the effectiveness of our user-controllable generative dehazing framework.
Appendix E Limitations
Although our method develops generative dehazing capabilities to handle dense haze and glow scenarios, its performance is inherently constrained by the pretrained diffusion model (e.g., Stable Diffusion). For instance, if a scene contains fine-grained elements such as text, signage, or uncommon objects that are underrepresented in the model’s training data, the generated content in heavily obscured regions may be less accurate or semantically inconsistent.
Appendix F Societal Impact
Nighttime visibility is often degraded by haze and glow effects, which limits the performance of applications such as autonomous driving and surveillance. Our method addresses these challenges by leveraging generative dehazing priors to restore or infer obscured content. This enhancement of nighttime visual quality can benefit downstream vision tasks and improve human interpretation in low-light, haze-affected environments.